OpenStudy (anonymous):

What is the slope-intercept form of the equation y-7 = -5/2 (x + 4)

OpenStudy (anonymous):

So is there mutiple choice

OpenStudy (anonymous):

Slope-intercept form of a line: y = mx + b y-7 = -5/2 (x + 4) y-7 = -5(x + 4) --------- 2 y= -5(x + 4) --------- +7 2 = -5x + -20 ----------- +7 2 = -5x - 10 + 7 ---- 2 = -5x ---- -3 2 y = -5/2x -3

OpenStudy (anonymous):

Okay